kittybasil:

QUERY: word problem? Not sure if this means the entire window or just the labeled portion

kittybasil:

I have the answer for a single and the full set of panels in the hexagon.

1 attachment
kittybasil:

I apologize for tagging two mods but you guys are the best math professionals I know out there @Vocaloid @angle

Angle:

What are you stuck with? Step 1: Solve for the area of 1 of the triangles using the given formula Step 2: Multiply by 6 because there are 6 triangles

kittybasil:

Well, I know how to solve the problem, except idk if it's asking for the entire window or not. Wording is unclear

Angle:

oh yeah entire window

kittybasil:

Especially since it says "windows" for some reason in the beginning. Extra confusion

Angle:

nono you're over thinking it the first sentence is saying that the sunlight for any window depends on the area of the glass of the window in general

Angle:

because it was talking about 'in general' it used plural in the first sentence

kittybasil:

Thank you so much! 🏅

kittybasil:

Protip for those who are studying/learning this kind of math:


To get a singular panel, use triangular area formula \(A=\frac{1}{2}bh\) where \(b=6\) and \(h=5.2\),\[\therefore A_{panel}=\frac{1}{2}(6)(5.2)=\frac{1}{2}{31.2}=15.6\text{ in}^2\]
To get the entire window, just multiply the previous value by 6 since there are six panels:\[A_{window}=6\cdot A_{panel}=6\cdot15.6\text{ in}^2=93.6\text{ in}^2\]
Hope that helped!

kittybasil:

I wish I could edit the original question to make it a reference post, but editing is bugged

kittybasil:

Actually, I had another question. Would this be \(0.77\overline{87}\) or \(0.7\overline{78}\)? (Since nobody knows if there's another 8 at the tail end)

1 attachment
Angle:

the latter 0.778 with the 78 repeating

Angle:

it gets highlighted for when the pattern first starts repeating

kittybasil:

I thought so too but idk if there's an 8 at the end after wait, highlighted?

Angle:

I meant with the line over it

kittybasil:

yeah, I wasn't sure which version it was due to not being sure if there is an 8 at the end (which there would be if it's \(0.7\overline{78}\) right?)

Angle:

there is no end

kittybasil:

oh I mean 8 at the end of revealed sequence, my bad

Angle:

it doesn't matter, the dot dot dots mean that whatever pattern you see should be repeated

kittybasil:

ah, okay but is \(0.77\overline{87}\) still valid though? That was my main issue because both look like they qualify as the answer, but there's only one 💀

kittybasil:

one correct answer*

Angle:

assume that mathematicians are lazy - always pick the most efficient notation

kittybasil:

😂 Okay

Angle:

it would be technically correct, but in terms of a multiple choice question - that is how I would think about it

kittybasil:

So then go with this - ?\[\large{0.7\overline{78}}\]

Angle:

correct

kittybasil:

Okay, ty!! double medals should be allowed - jk

Angle:

well you see... this is why we only allow one question per post

kittybasil:

oh no I was joking haha

kittybasil:

Recap for those who are trying to sort through the thread. Here is a summary of all the content discussed!

kittybasil wrote:
Protip for those who are studying/learning this kind of math: To get a singular panel, use triangular area formula \(A=\frac{1}{2}bh\) where \(b=6\) and \(h=5.2\),\[\therefore A_{panel}=\frac{1}{2}(6)(5.2)=\frac{1}{2}{31.2}=15.6\text{ in}^2\] To get the entire window, just multiply the previous value by 6 since there are six panels:\[A_{window}=6\cdot A_{panel}=6\cdot15.6\text{ in}^2=93.6\text{ in}^2\]
Also, recap of the second question: \(0.778787...\) should translate to this:\[\large{0.7\overline{78}}\]